# Quantum information with quantum-like bits

### v4

@@ IV.2 Quantum probabilities in a QL model
$=\sum_{j=1}^{J}P^{(\mathrm{cl})}\left(\mathcal{B}\cap A_{j}\right)=\sum_{j=1}^{J}P^{(\mathrm{cl})}(\mathcal{B}|\mathcal{A}_{j})P^{(\mathrm{cl})}(\mathcal{A}_{j}),$ (IV.7)
@@
$P^{(\mathrm{cl})}(\mathcal{B}|\mathcal{A}_{j})=\frac{P^{(\mathrm{cl})}(\mathcal{B}\cap\mathcal{A}_{j})}{P^{(\mathrm{cl})}(\mathcal{A}_{j})}.$ (IV.8)
We discuss now the correspondent of ?? for the present QL framework. Let $\mathcal{F}$ and $\mathcal{G}$ be two nondegenerate QL observbles, whose eigenvalue problems are defined as in ??. In this case, the QL conditional probabilities defined in ?? lose their dependence on the initial state $\psi$:
$P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)$ $\mapsto|\braket{g_{\bm{\sigma}^{\prime}},f_{\bm{\sigma}}}|^{2}\equiv P\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right).$ (IV.9)
@@
$P\left(\lambda^{\mathcal{G}}_{\bm{\sigma}}|\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime}}\right)=\left.P_{\psi}(\lambda^{\mathcal{G}}_{\bm{\sigma}})\right|_{\psi=f_{\bm{\sigma}^{\prime}}}.$ (IV.10)
@@
×$\quad\times P_{\psi}\left(\lambda_{\bm{\sigma}^{\prime\prime}}^{\mathcal{F}}\right)P\left(\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime\prime}}|\lambda^{\mathcal{G}}_{\bm{\sigma}}\right)\Big{]}^{1/2}.$ (IV.11)
@@
V End-to-End quantum-like information processing and measurement
For genuine quantum resources, the primary source of error during information processing and measurement arises from the presence of uncontrolled noise. This can result from a variety of sources, including external electromagnetic fields, undesired interactions between qubits, fabrication defects, and imperfections in control protocols. Hou et al. (2016); Bengtsson et al. (2024); Funcke et al. (2022); Gupta and Nene (2021); Resch and Karpuzcu (2021) To account for these sources of error the final state of a quantum register after computation requires multiple measurements, which can lead to a significant overhead in the computing protocols. Likewise, error mitigation techniques and decoding protocols can significantly affect scalability and performance. Bravyi et al. (2021); Ali et al. (2024). In the following, we analyze both the accuracy and computational complexity of information processing and measurement within the QL formalism.
Given that QL resources are intrinsically classical, noise from external sources of dissipation does not appear to be a fundamental limitation in our approach. Instead, the sources of errors are intrinsic to the formalism itself, impacting on an equal footing information processing and final measurement. Here, we present an analysis of these errors by examining the overall workflow of QL information protocols from end to end.
Let us consider a generic QL algorithm consisting of a set of $M$ gates $\mathrm{g}=\{g_{1},\cdots,g_{M}\}$. This corresponds to the chain of transformations $U_{\bm{\mathrm{g}}}=\prod_{j=M}^{1}U_{g_{j}}$, as defined in ??. An exact approach to QL computing would involve applying this sequence of gates to transform via ?? the initial register $\mathcal{R}$ [??]. This consists of the Cartesian product of ${N_{\mathrm{QL}}}$ factors, as illustrated in panel a) of ??. Hence, storing and processing information following this exact strategy involves an exponential overhead with the number of edges in the total resource. A key insight we aim to leverage to reduce the complexity of this approach is as follows: if we can develop approximate QL protocols that maintain the structure of the Cartesian product, the exponential operations required for information processing and measurement could be circumvented entirely. We refer to this approach as the Cartesian-product approximation (CPA). This strategy enables to implement gate operations by transforming only a small subset of the edges of the total resources. In fact, within the CPA, information can be conveniently stored and processed with linear complexity in ${N_{\mathrm{QL}}}$ [see panel b) of ??].
Firstly, we assess the degree to which the CPA holds within our approach and quantify the associated errors relative to exact QL computing. Conveniently, the Cartesian-product structure is exactly preserved by all one-QL-bit gates , defined by
$W^{(p)}_{\mathrm{g}}=\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes U^{(p)}_{\mathrm{g}}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)},$ (V.1)
acting locally on a single resource $p\in\{1,\cdots,{N_{\mathrm{QL}}}\}$. ?? transforms the initial graph resource ?? as
$\displaylimits_{q=1}^{N_{\mathrm{QL}}}\mathcal{G}^{(q)}\;\begin{subarray}{c}U^{(p)}_{\mathrm{g}}\\ \longmapsto\end{subarray}\;\mathcal{G}^{(1)}\square\cdots\square\mathcal{G}^{(p)}_{\mathrm{g}}\square\cdots\square\mathcal{G}^{({N_{\mathrm{QL}}})},$ (V.2)
where $\mathcal{G}^{(p)}_{\mathrm{g}}$ is the gate transformation of the graph $\mathcal{G}^{(p)}$. To extend the approach to a universal set of gates, we analyze how effectively the CPA approximation holds for the CNOT operation $U_{\mathrm{CNOT}}^{(q,p)}$, applying the Pauli-X gate to the $p$-th QL bit depending on the state of the control $q$. As shown in ??, this gate conserves the Cartesian product only approximately, namely
$\displaylimits_{q=1}^{N_{\mathrm{QL}}}\mathcal{G}^{(q)}\;$ $\begin{subarray}{c}U_{\mathrm{CNOT}}^{(q,p)}\\ \longmapsto\end{subarray}\;\mathcal{G}^{(1)}\square\cdots\square\tilde{\mathcal{G}}^{(p)}\square\cdots\square\tilde{\mathcal{G}}^{(q)}\square\cdots\square\mathcal{G}^{({N_{\mathrm{QL}}})}$
$\quad+\mathcal{O}(l/k).$ (V.3)
?? introduces an error which depends on the ratio between the valencies $l$ and $k$ of the elementary subgraphs in ??, respectively $C$ and $A_{i}$. This result indicates that the optimal choice for the parameters defining the subgraphs are
$l=1,\hskip 14.22636ptk={N_{G}}-1,\hskip 14.22636pt{N_{G}}\gg 1.$ (V.4)
In future work, we aim to investigate the propagation of error in ?? as several gates are concatenated to construct QL algorithms.
An illustration of the final measurement process after information processing is shown in panel c) of ??. Let us firstly consider factorized measurement operators, with structure
$\mathcal{F}_{{\bm{g}}^{\prime}}=\bigotimes_{q=1}^{N_{\mathrm{QL}}}U^{(q)}_{g^{\prime(q)}}.$ (V.5)
Conveniently, even at the measurement state we can avoid to explicitly calculate the exponentially expensive Cartesian product for this common class of observable operators. In fact, as shown in panel c) of ??, local unitary transformations applied to one-QL-bit resources provide complete information on the measurement probabilities ??. For example, a local measurement of the population difference of the $q$-th QL bit acts on the final local emergent states as
$\psi^{(q)}_{\sigma^{(q)},\bm{\mathrm{g}}}$ $\mapsto U_{\mathrm{Z}}^{(q)}\psi^{(q)}_{\sigma^{(q)},\bm{\mathrm{g}}},$ (V.6a)
[…]
The more general case of non-factorized observables involves the sum over $M$ terms ?? expressed by
$\mathcal{F}_{\bm{g}_{1}^{\prime},\cdots,\bm{g}_{M}^{\prime}}=\sum_{m=1}^{M}\mathcal{F}_{\bm{g}^{\prime}_{m}}.$ (V.7)
The action of ?? breaks the structure of the Cartesian product, leading to a measurement cost that scales exponentially with the number of QL bits. It is worth noting that in a broad class of quantum algorithms, including Grover’s search Giri and Korepin (2017) and variational quantum eigensolvers, Tilly et al. (2022) the measurement process involves exclusively factorized operators, with structure as in ??. In these cases, this local measurement process can be applied with no limitations. We aim to compile a comprehensive list of the potential applications of our approach in future work.
Figure 5: a) Representation of the initial register, prior to information processing. The number of vertices of the total graph scales exponentially with ${N_{\mathrm{QL}}}$, the number of QL-bits. b) Information processing implemented within the CPA, as discussed in the main text. Here, we store and transform the single resources individually, with a favorable linear scaling with size. c) Illustration of the final measurement process. This involves the local diagonalization of the gate-transformed resources, to obtain ${N_{\mathrm{QL}}}$ sets of single-particle emergent states. Factorized measurements operators can be implemented locally on these resources, without calculating the tensor product associated to the final global emergent state.
VI Conclusions
@@
In this paper, we expanded upon the general approach derived in Refs. Scholes, 2023, 2024; Scholes and Amati, 2024 to define and manipulate QL states. In ?? we constructed a computational basis from the emergent states of many-body classical systems, whose correlations can be mapped onto synchronizing graphs. We introduced the Hilbert space for an arbitrary number of QL bits, and we derived explicitly the Bell states in dimension two. In ?? we proposed a notion of quantum gates applicable to QL states, and we provided concrete examples of their action. We discussed in ?? a list of measurement axioms which allows to observe QL interference effects in our approach. Finally, we outlined in ?? an efficient end-to-end workflow for building QL algorithms based on a controlled approximation for two-QL-bit gates.
@@
In the paper we introduced notions of entanglement, interference, and QL probabilities which avoid the expensive calculation of joint probability distributions in the measurement process. In future work we plan to investigate how each of these features contributes to achieving a “QL advantage” over strictly classical information processing. The ultimate test of the approach proposed in this work will be to implement the pipeline proposed in ?? (and related ??) to study QL algorithms, and evaluate their time efficiency and resource demand. Bharti et al. (2022)
@@
A broad class of computational approaches to open quantum systems and quantum chemistry involves a perspective somehow complementary to our approach. In particular, quantum nonadiabatic systems are often mapped onto quasiclassical models to minimize computational complexity and inexpensively approximate long-time dynamics. Accurate quasiclassical methods have been derived, by requiring that Kolmogorov probabilities [fulfilling ??] obey by construction quantum detailed balance at long times. Amati, Mannouch, and Richardson (2023); Amati, Runeson, and Richardson (2023); Runeson et al. (2022); Amati et al. (2022); Mannouch and Richardson (2023) It can be insightful to assess whether even more accurate quasiclassical approaches can be derived from a QL stochastic theory which approximates ?? beyond the first classical term.
In this paper, we introduced a universal set of gates comprising single QL-bit terms and the CNOT operation. Our controlled approximation is designed to preserve the Cartesian product structure by selectively neglecting a small contribution within the CNOT gate that would otherwise lead to symmetry breaking. Preserving this contribution would result in an exponential increase in complexity, posing significant challenges for both computational and experimental resources. An open question remains as to whether directly implementing an approximation of many-QL-bit gates, rather than decomposing them into one- and two-body elementary terms, could yield even greater accuracy. While any local decomposition is exact in standard unitary quantum computing, this is in general not the case within our QL approach. Additionally, alternative choices of universal sets may provide greater computational advantages. In future work, we aim to explore the impact of these different scenarios within our framework.
From an experimental standpoint, various possibilities can be considered for implementing synchronizing resources in QL information processing. The research on oscillator-based computing, as reviewed in Ref. Csaba and Porod, 2020, offers a solid foundation for this purpose. Within this framework, insulator-to-metal-transition devices and spin-torque oscillators emerge as promising candidates, supporting distributed and parallel computing while ensuring resilience to noise and fault tolerance.Raychowdhury et al. (2019) We plan to further investigate platforms for the experimental implementation of QL resources, focusing on the properties needed by various approaches to achieve optimal network design and control.
In recent years, memcomputing has emerged as an innovative computational framework that exploits the intrinsic time-nonlocal properties of classical dynamical systems to enable a terminal-agnostic framework to implement non-Boolean logic. Di Ventra and Traversa (2018) This approach, which allows the computation to occur in memory rather than through sequential logic operations, could be exploited by recasting classical synchronizing dynamics leading to emergent states into time-nonlocal equations of motion. Amati, Meyer, and Schilling (2019) The classical Nakajima–Mori–Zwanzig formalism, which captures the influence of past states on open classical systems, offers a strategy to incorporate memory effects directly into the QL formalism, potentially enhancing stability and coherence.
The environmental implications of the large-scale industrial production of quantum processors are topics of ongoing debate. Ezratty (2023); Ikonen, Salmilehto, and Möttönen (2017); Paler and Basmadjian (2022) For instance, superconducting devices require significant energy due to the cryogenic cooling of the constituent superconducting circuits. Similarly, quantum computers based on arrays of Rydberg atoms demand intensive energy resources. Meyer (2000); Bernien et al. (2017) Implementing energy-efficient classical circuits for QL computing can contribute to sustainable quantum-processing engineering. Additionally, our method leverages classical synchronization, which scales more efficiently with resources compared to its quantum-mechanical counterpart. Aifer, Thingna, and Deffner (2024)
@@
VII Acknowledgments
This research was funded by the National Science Foundation under Grant No. 2211326 and the Gordon and Betty Moore Foundation through Grant GBMF7114.
Appendix A Transformation of the emergent states to the computational basis
@@ Appendix B Derivation of the quantum-like formula of total probabilities
$u_{\bm{\sigma}^{\prime},\bm{\sigma}}$ $=\braket{g_{\bm{\sigma}},f_{\bm{\sigma}^{\prime}}}.$ (B.2b)
@@ Appendix C Spectral properties of QL observables
Appendix D Controlled-NOT gate
In this appendix, we analyze the action of the CNOT gate on a pair of QL bits, $p,q\in{1,\dots,{N_{\mathrm{QL}}}}$. Here, we designate $p$ as the control QL bit, whose state determines whether to flip the state of the target QL bit $q$. For simplicity, we assume $p<q$; the case $p>q$ can be derived analogously. The unitary map implementing the transformation is
$U_{\mathrm{CNOT}}^{(p,q)}$ $=\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\uparrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}$
$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\downarrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes U_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)},$ (D.1)
where we defined the projection operators
@@ Appendix D Controlled-NOT gate
$\mathcal{P}_{i}^{(p)}=\tfrac{1}{2}\left[U_{\mathcal{I}}^{(p)}+(-1)^{i}U_{\mathrm{Z}}^{(p)}\right],$ (D.2)
with $p\in\{1,\cdots,{N_{\mathrm{QL}}}\}$ and $i\in\{0,1\}$. These are mutually orthogonal [$\mathcal{P}_{\sigma}^{(p)}\mathcal{P}_{\sigma^{\prime}}^{(p)}=\delta_{\sigma\sigma^{\prime}}\mathcal{P}_{\sigma^{\prime}}^{(p)}$] and complementary [$\mathcal{P}_{\downarrow}^{(p)}+\mathcal{P}_{\uparrow}^{(p)}=U_{\mathcal{I}}^{(p)}$]. The action of ?? on the total resource ?? is
$\mathcal{R}_{\mathrm{CNOT}(q,p)}^{[{N_{\mathrm{QL}}}]}$ $=\sum_{\sigma\in\{\uparrow,\downarrow\}}\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{R}_{\sigma\sigma}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}+\sum_{\begin{subarray}{c}r=1\\ r\neq q\end{subarray}}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}^{(r)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}$ (D.3a)
$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\uparrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes\mathcal{R}^{(q)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-q)}+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{P}_{\downarrow}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes\mathcal{R}_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)}$ (D.3b)
$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\left[\mathcal{R}_{\uparrow\downarrow}^{(p)}+\mathcal{R}_{\downarrow\uparrow}^{(p)}\right]\otimes\mathbb{1}_{2{N_{G}}}^{\otimes(q-p-1)}\otimes U_{\mathrm{X}}^{(q)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-q)},$ (D.3c)
where $\mathcal{R}^{(p)}_{\sigma\pi}=\mathcal{P}^{(p)}_{\sigma}\mathcal{R}^{(p)}\left[\mathcal{P}^{(p)}_{\pi}\right]^{\dagger}$. As introduced in ??, we can minimize the computational complexity of this gate by assessing to which extent the structure of ?? can be approximated by a Cartesian product. This property enables information through transformations applied exclusively to individual one-QL-bit resources. Let us therefore study separately all terms which would break the Cartesian product construction. Starting from the two terms in ??, we can study the relative error
$\frac{\left\lVert\mathcal{R}_{\mathrm{X}}^{(q)}-\mathcal{R}^{(q)}\right\rVert}{\left\lVert\mathcal{R}^{(q)}\right\rVert}=\sqrt{\frac{8l{N_{G}}}{2k{N_{G}}+2l{N_{G}}}}=\frac{2}{\sqrt{1+\frac{k}{2l}}}\simeq 1+\frac{2l}{k}.$ (D.4)
In ?? we introduced the Frobenius norm, defined by $\left\lVert A\right\rVert^{2}=\sum_{i=1}^{m}\sum_{j=1}^{n}|A_{ij}|^{2}$, for each $A\in M_{m\times n}(\mathbb{C})$. Also, we noticed that
$\mathcal{R}^{(q)}_{\mathrm{X}}=U_{\mathrm{X}}^{[q]}\mathcal{R}^{(q)}\left[U_{\mathrm{X}}^{[q]}\right]^{\dagger}=\begin{pmatrix}A_{1}^{(q)}&-C^{(q)}\\ -\left[C^{(q)}\right]^{\top}&A_{2}^{(q)}\end{pmatrix}.$ (D.5)
?? is minimized by the parameter choice defined in ??. Finally, let us consider the cross contribution in ??. This terms involves the difference
$\mathcal{R}^{(p)}_{\uparrow\downarrow}+\mathcal{R}^{(p)}_{\downarrow\uparrow}$ $=\frac{1}{2}(\mathcal{R}^{(p)}-\mathcal{R}^{(p)}_{Z})$
$=\frac{1}{2}\begin{pmatrix}A_{1}^{(p)}-A_{2}^{(p)}&C^{(p)}-\left[C^{(p)}\right]^{\top}\\ \left[C^{(p)}\right]^{\top}-C^{(p)}&A_{2}^{(p)}-A_{1}^{(p)}\end{pmatrix},$ (D.6)
where
$\mathcal{R}^{(q)}_{\mathrm{Z}}=\begin{pmatrix}A_{2}^{(q)}&\left[C^{(q)}\right]^{\top}\\ C^{(q)}&A_{1}^{(q)}\end{pmatrix}$ (D.7)
is calculated similarly to ??. Without loss of generality it is possible to choose here
$A_{1}^{(p)}$ $=A_{2}^{(p)},$ (D.8a)
$C^{(p)}$ $=\left[C^{(p)}\right]^{\top},\hskip 14.22636pt\forall\;p=1,\cdots,{N_{\mathrm{QL}}}.$ (D.8b)
In essence, we obtain
$\mathcal{R}_{\mathrm{CNOT}(q,p)}^{[{N_{\mathrm{QL}}}]}$ $=\sum_{\sigma\in\mathcal{S}}\mathbb{1}_{2{N_{G}}}^{\otimes(p-1)}\otimes\mathcal{R}_{\sigma\sigma}^{(p)}\otimes\mathbb{1}_{2{N_{G}}}^{\otimes({N_{\mathrm{QL}}}-p)}$
$\quad+\sum_{\begin{subarray}{c}r=1\\ r\neq q\end{subarray}}^{{N_{\mathrm{QL}}}}\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}^{(r)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}$
$\quad+\mathbb{1}_{2{N_{G}}}^{\otimes(r-1)}\otimes\mathcal{R}_{\mathrm{X}}^{(q)}\otimes\mathbb{1}^{\otimes({N_{\mathrm{QL}}}-r)}+\mathcal{O}(l/k)$ (D.9)
provided the conditions in ?? hold.

### v5

@@ VI.2 Quantum probabilities in a QL model
$=\sum_{j=1}^{J}P^{(\mathrm{cl})}\left(\mathcal{B}\cap A_{j}\right)=\sum_{j=1}^{J}P^{(\mathrm{cl})}(\mathcal{B}|\mathcal{A}_{j})P^{(\mathrm{cl})}(\mathcal{A}_{j}),$ (VI.7)
@@
$P^{(\mathrm{cl})}(\mathcal{B}|\mathcal{A}_{j})=\frac{P^{(\mathrm{cl})}(\mathcal{B}\cap\mathcal{A}_{j})}{P^{(\mathrm{cl})}(\mathcal{A}_{j})}.$ (VI.8)
We discuss now the correspondent of ?? for the present QL framework. We restrict the following analysis to couples of non-degenerate QL observables, $\mathcal{F}$ and $\mathcal{G}$, whose eigenvalue problems are defined as in ??. The eigenvalues of these operators are associated with a one-dimensional subspaces, which can be fixed by selecting any normalized eigenvector. In this case, the QL conditional probabilities defined in ?? lose their dependence on the initial state $\psi$, and can be simply expressed as
$P_{\psi}\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right)$ $\mapsto|\braket{g_{\bm{\sigma}^{\prime}},f_{\bm{\sigma}}}|^{2}\equiv P\left(\lambda^{\mathcal{G}}_{\bm{\sigma}^{\prime}}|\lambda^{\mathcal{F}}_{\bm{\sigma}}\right).$ (VI.9)
@@
$P\left(\lambda^{\mathcal{G}}_{\bm{\sigma}}|\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime}}\right)=\left.P_{\psi}(\lambda^{\mathcal{G}}_{\bm{\sigma}})\right|_{\psi=f_{\bm{\sigma}^{\prime}}}.$ (VI.10)
@@
×$\quad\times P_{\psi}\left(\lambda_{\bm{\sigma}^{\prime\prime}}^{\mathcal{F}}\right)P\left(\lambda^{\mathcal{F}}_{\bm{\sigma}^{\prime\prime}}|\lambda^{\mathcal{G}}_{\bm{\sigma}}\right)\Big{]}^{1/2}.$ (VI.11)
@@
VII Conclusions
@@
In this paper, we built upon the general framework developed in Refs. Scholes, 2023, 2024; Scholes and Amati, 2025 to define and manipulate QL states. In ??, we constructed a computational basis from the emergent states of many-body classical systems, whose correlations can be mapped onto synchronizing graphs. We introduced the Hilbert space for an arbitrary number of QL bits and explicitly derived the Bell states in two dimensions. In ??, we proposed a framework for defining quantum gates applicable to QL states and provided concrete examples of their action on many-body classical resources. ?? offers an overview of experimental platforms that could be used to implement and validate QL theory. Finally, in ??, we discussed that a QL notion of interference can emerge within our framework by presenting a set of axioms for interpreting QL states and observables.
@@
A broad class of computational approaches to open quantum systems and quantum chemistry involves a perspective somehow complementary to our approach. In particular, quantum nonadiabatic systems are often mapped onto quasiclassical models to minimize computational complexity and inexpensively approximate long-time dynamics. Accurate quasiclassical methods have been derived, by requiring that Kolmogorov probabilities [fulfilling ??] obey by construction quantum detailed balance at long times. Amati, Mannouch, and Richardson (2023); Amati, Runeson, and Richardson (2023); Runeson et al. (2022); Amati et al. (2022); Mannouch and Richardson (2023); Amati (2024) It can be insightful to assess whether even more accurate quasiclassical approaches can be derived from a QL stochastic theory which approximates ?? beyond the first classical term.
In this paper, we presented an early-stage analysis exploring the connection between QL gates and the many-body synchronization dynamics of oscillator systems. Specifically, we identified relationships between the structure of the adjacency matrix of the network and the synchronization patterns observed in the Kuramoto model. For future work, we aim to develop a more formal framework by deriving general expressions for multi-time correlation functions associated with network transformations induced by chains of QL gates. Additionally, we plan to investigate whether the structure of emergent QL states can be directly inferred from the synchronizing steady states and how this process is influenced by the microscopic details of the underlying classical dynamics. Future work could explore new possibilities regarding the generality of the present approach for different synchronizing dynamical systems. While this study focused on the nonlinear interactions of the Kuramoto model, it is crucial to assess the applicability of the results to a broader class of classical systems exhibiting synchronization. Extending this framework to more general models, where oscillators are represented by vectorial variables de Aguiar (2023) or elements of a unitary group, as in the Lohe model Lohe (2009); Scholes (2022), offers a promising path for further research. These generalizations enable the encoding of high-dimensional information within the phase relationships between graph nodes, potentially enhancing the modeling of complex dynamical systems.
In recent years, memcomputing has emerged as a new computational paradigm that leverages the intrinsic time-nonlocal properties of nonlinear classical dynamical systems to enable non-Boolean logic computations by exploiting memory effects within physical systems. A key feature of memcomputing is its terminal-agnostic nature, where input and output terminals are treated equivalently. This allows self-organizing circuits to dynamically adjust and converge to equilibrium points that correspond to solutions, regardless of the terminal through which data is introduced or extracted. Di Ventra and Traversa (2018) The feasibility of the memcomputing framework has been demonstrated across various physical implementations, including oscillator-based systems. Traversa et al. (2015); Di Ventra (2022) It can be insightful to implement a variant of our approach with flavors of memcomputing techniques, by exploiting time-nonlocality dissipative nonlinear dynamical systems, e.g. as predicted within the Nakajima–Mori–Zwanzig formalism. Nakajima (1958); Mori (1965); Zwanzig (1960); Dominy and Venturi (2017); Amati et al. (2022); Amati, Meyer, and Schilling (2019)
The environmental implications of large-scale industrial production of quantum processors remain a topic of ongoing debate. Ezratty (2023); Ikonen, Salmilehto, and Möttönen (2017); Paler and Basmadjian (2022); Sood and Chauhan (2024); Jaschke and Montangero (2023); Auffèves (2022) Conducting an early-stage assessment of the environmental impact of various resources can provide valuable metrics to identify the most energy-efficient and scalable quantum technologies for future development. In this context, implementing energy-efficient classical circuits for QL computing offers a promising pathway toward sustainable QL-processing engineering. While our perspective does not claim to provide a definitive solution, we recognize that classical systems, despite their larger size, could serve as viable energy-efficient alternatives for specific use cases, such as small-scale, proof-of-concept analyses in fundamental research.
@@
VIII Acknowledgments
We thank Debadrita Saha for useful discussions. This research was funded by the National Science Foundation under Grant No. 2211326 and the Gordon and Betty Moore Foundation through Grant GBMF7114.
Appendix A General transformation from emergent states to the computational basis
@@ Appendix B Derivation of the quantum-like formula of total probabilities
$u_{\bm{\sigma}^{\prime},\bm{\sigma}}$ $=\braket{g_{\bm{\sigma}},f_{\bm{\sigma}^{\prime}}},$ (B.2b)
@@ Appendix C Spectral properties of QL observables
Appendix D Dynamics of the Kuramoto model
Here, we present a closed-form solution for the dynamics of the Kuramoto model, as described by the nonlinear equations of motion in ??. Our approach follows the methods outlined in Refs. Muller, Mináč, and Nguyen, 2021; Kawamura et al., 2010. This analysis offers a practical framework for implementing the dynamics of the model efficiently, without the need for explicitly integrating the equations of motion, making it applicable to a broad range of synchronizing networks. Additionally, this work can serve as a starting point for future analytical studies exploring the role of emergent states in synchronization processes.
If the coupling matrix $\mathcal{R}$ is real, it is convenient to subtract from ?? an additional complex component,
$\dot{\theta}_{i}$ $=\omega_{i}+\frac{\Gamma}{2{N_{G}}}\sum_{j=1}^{2{N_{G}}}\mathcal{R}_{ij}\left[\sin(\theta_{j}-\theta_{i})-i\cos(\theta_{j}-\theta_{i})\right],$ (D.1)
leading to
$i\dot{\theta}_{i}=i\omega_{i}+\frac{\Gamma}{2{N_{G}}}\,\mathrm{e}^{-i\theta_{i}}\sum_{j=1}^{N_{\mathrm{QL}}}\mathcal{R}_{ij}\,\mathrm{e}^{i\theta_{j}}.$ (D.2)
?? can be expressed in matrix form as
@@ Appendix D Dynamics of the Kuramoto model
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathrm{e}^{i\bm{\theta}}=\left[{\mathrm{diag}}(i\bm{\omega})+\frac{\Gamma}{2{N_{G}}}\mathcal{R}\right]\,\mathrm{e}^{i\bm{\theta}}$ (D.3)
or, more compactly as, $\dot{\bm{x}}=\tilde{\mathcal{R}}\bm{x}$, where $\bm{x}=\,\mathrm{e}^{i\bm{\theta}}=(\,\mathrm{e}^{i\theta_{1}},\cdots,\,\mathrm{e}^{i\theta_{2{N_{G}}}})^{\top}$ and ~$\tilde{\mathcal{R}}=\mathrm{diag}(i\bm{\omega})+\frac{\Gamma}{2{N_{G}}}\mathcal{R}$. This allows us to obtain as a closed form for the solution of the dynamics of the Kuramoto model
$\bm{x}(t)=\,\mathrm{e}^{\tilde{\mathcal{R}}t}\bm{x}(0).$ (D.4)
We can decompose the real and imaginary part of the frequency vector $\bm{\theta}=\bm{\theta}_{\mathrm{re}}+i\bm{\theta}_{\mathrm{im}}$, to obtain as a physical solution of ??
$i\bm{\theta}_{\mathrm{re}}(t)-\bm{\theta}_{\mathrm{im}}(t)$ $=\log\left[\,\mathrm{e}^{\tilde{\mathcal{R}}t}\bm{x}(0)\right],$ (D.5)
or
$\bm{\theta}_{\mathrm{re}}(t)$ $=\mathrm{Im}\left\{\log\left[\,\mathrm{e}^{\tilde{\mathcal{R}}t}\bm{x}(0)\right]\right\},$ (D.6a)
$\bm{\theta}_{\mathrm{im}}(t)$ $=-\mathrm{Re}\left\{\log\left[\,\mathrm{e}^{\tilde{\mathcal{R}}t}\bm{x}(0)\right]\right\}.$ (D.6b)